OpenStudy (moneybaby980):

A proportional relationship between the number of pounds of cabbage (x) and the price in dollars (y) is graphed, and the ordered pair (5, 2) is on the graphed line. Part A: What is the price of 1 pound of cabbage? Show your work, including the proportion you used to determine the price. Part B: What does the ordered pair (10, 4) on the graph represent? Explain in words.

OpenStudy (anonymous):

I had a question like this but the problem was (5,4) not (5,2)I did this, Part A: What is the price of 1 pound of potatoes? Show your work, including the proportion you used to determine the price. (8 points) Part B: What does the ordered pair (10, 8) on the graph represent? Explain in words. the relationship is linear in the form y = mx +c If (5,4) is on the line we can calculate the equation gradient m = change in y over change in x, = 4/5 y = 4/5 * x +c 4 = 4/5 * 5 +c c = 0 y = 4/5 x Part a if x = 1 then y = 4/5 part b is 2 times (5,4) so 10 pounds of potatoes cost 8 dollars

OpenStudy (moneybaby980):

were did u get two from